Solve for x: 4x^2-21x^2+27=0
Answer by mananth(16946)

4x^2-21x^2+27=0
..
-17x^2+27=0
-17x^2= -27
x^2= -27/-17
x^2 = 27/17
x= + / - sqrt (27/17)